When an aqueous solution containing gold (III) salt is electrolyzed, metallic


gold is deposited at the cathode and oxygen gas is generated at the anode. If


9.26 g of Au is deposited at the cathode, calculate the volume (in liters) of O2


generated at 23°C and 747 mmHg.

1
Expert's answer
2021-12-21T20:32:01-0500

Mole ratio for discharge is:

Au : O

2 : 3.

Moles of oxygen discharged was

3/2 × 9.26 = 13.89 moles


Volume of oxygen :

At RTP, molar gas volume is 24dm3

Volume of 13.89 moles of oxygen is

13.89×24=333.36litres


P1V1/T1=P2V2/T2

296×747/V2 = 298×760/333.36

And = 325.46 litres